# Efficient detection of chaos through the computation of the Generalized Alignment Index (GALI) by the multi-particle method

### v1

@@ Abstract
We present a thorough analysis of computing the Generalized Alignment Index (GALI), a rapid and effective chaos indicator, through a simple multi-particle approach, which avoids the use of variational equations. We develop a theoretical leading-order estimation of the error in the computed GALI for both the variational method (VM) and the multi-particle method (MPM), and confirm its predictions through extensive numerical simulations of two well-known Hamiltonian models: the Hénon-Heiles and the $\beta$-Fermi-Pasta-Ulam-Tsingou systems. For these models the GALIs of several orders are computed and the MPM results are compared to the VM outcomes. The dependence of the accuracy of the MPM on the renormalization time, integration time step, as well as the deviation vector size, is studied in detail. We find that the implementation if the MPM in double machine precision ($\varepsilon\approx 10^{-16}$) is reliable for deviation vector magnitudes centred around $d_{0}\approx\varepsilon^{1/2}$, renormalization times $\tau\lesssim 1$, and relative energy errors $E_{r}\lesssim\varepsilon^{1/2}$. These results are valid for systems with many degrees of freedom and for several orders of the GALIs, with the MPM particularly capturing very accurately the $\textrm{GALI}_{2}$ behavior. Our results show that the computation of the GALIs by the MPM is a robust and efficient method for investigating the global chaotic dynamics of autonomous Hamiltonian systems, something which is of distinct importance in cases where it is difficult to explicitly write the system’s variational equation or when these equations are too cumbersome.
@@ 1 Introduction
Revealing the properties of individual or ensembles of orbits belonging to particular phase space region is of fundamental interest for nonlinear systems. One of the most prominent tools for this task are chaos indicators, which are real-valued quantities exhibiting distinct behaviors for different dynamical evolutions [1, 2, 3, 4].
Over the last fifty years, several chaos detection techniques have been developed and employed in a wide variety of fields. The Lyapunov exponents (LEs) [5, 6] for example, have become a well-established cornerstone in understanding and quantifying chaos in nonlinear systems (see e.g. [7, 3, 8] and references therein). Alongside with them, other chaos detection techniques, such as the fast Lyapunov indicator (FLI) [9, 10], the mean exponential growth of nearby orbits (MEGNO) [11, 12] the smaller (SALI) and generalized (GALI) alignment indices [13, 14, 15], to name a few, have found applications in information theory [16], biophysics and materials [17, 18, 19], astrophysics [20, 21, 22], as well as population dynamics and epidemiology [23, 24]. In these works, the utility of each indicator in a different context was shown, emphasizing the efficiency with which the accurate chaos detection was achieved [14, 22]. This wide-ranging applicability of chaos indicators provides strong motivation for probing their evaluation, whether through optimising the numerical procedures [25, 26, 27], choosing appropriate methods of calculation [28, 29], or using short time computations to determine chaoticity [30]. Being able to reliably and efficiently evaluate these indicators in a variety of systems is of significant importance.
The majority of these chaos indicators rely on the time evolution of initially neighboring orbits to the trajectory under study. Depending on whether the neighboring orbits are kept in the vicinity or are driven away from the system’s reference orbit allows the characterization of the studied orbit’s long-time stability. In order to numerically evolve these initially nearby orbits, two methods are usually implemented. In the first case, alongside the system’s reference orbit, several initially nearby trajectories are evolved by numerically solving the system’s equations of motion. This is done in such a way that the time evolution of the separation vectors between the studied orbit and its neighbors can be tracked for all time, so that these vectors can be used to estimate the spatiotemporal properties of chaos indicators. A second approach is based on the fact that the initial separation vectors are by definition considered to be very small. Consequently their dynamics can be approximated by a linearization of the system’s equations of motion along the studied orbit. The equations obtained by this approximation are referred to as variational equations. In this work we name the former approach the multi-particle method (MPM) and the latter the variational method (VM).
Nowadays the VM is the most commonly used method, mainly due to its accuracy and simplicity. This approach relies on the explicit derivation of the variational equations, which requires the dynamical system’s force field to be continuous and differentiable over the considered space-time domain, something which is not always the case. As an example we can mention a model whose particles’ interactions are described via Hertzian contacts [31], or simply consider discontinuous or non-analytic force fields related to work softening [32, 33] and hardening [34, 35] laws. Furthermore, the ever increasing complexity of the equations describing the system’s dynamics in modern physical models makes the derivations of the variational equations difficult (see e.g. [19]). Thus, occasionally, even in cases where the variational equations can be explicitly expressed in closed analytical forms, their complexity could result to strains during the code implementation due to difficulties with respect to the code optimization (because of its large size), and due to the increase of computational times, as complex variational equations can considerably increase the number of basic computer operations.
The MPM can be implemented in a variety of cases, and seems to be particularly advantageous whenever the VM’s use becomes problematic, since it requires only the knowledge of the system’s equations of motion. Nevertheless, the results obtained from the MPM must always be treated with care because they are known to lead to spurious estimates, see Ref. [28] for examples in astrophysics. Consequently several studies of the reliability of the MPM for the computation of the MLE have already been carried out [28, 26, 29] (see also Sect. 3.5 of [3]). On the other hand, some more recently introduced chaos indicators such as the GALI method, have only been used in the framework of the VM. It follows that its definitions, reliability and computational performances in the context of the MPM has, to the best of our knowledge, never been examined.
@@
The aim of our study is to investigate the reliability of the GALI computation through the MPM. We start our analysis by demonstrating that the definitions of the GALI using the VM (see e.g. [14]) and the MPM are equivalent. Then, we obtain leading order analytical expressions of the numerical accuracy during the computation of the GALI values for both methods. We find that our analytical results are validated by simulating the short-time dynamics of some well-known, prototypical nonlinear Hamiltonian systems: the Hénon-Heiles (HH) [36] and $\beta-$Fermi-Pasta-Ulam-Tsingou ($\beta-$FPUT) [37, 38] systems. Moreover, using the parameter values which we found to optimize the performance of the MPM, we show that the GALI method is in general able to characterize with good accuracy the nature of the orbits of the studied systems.
We also examine the ability of the GALI method to reveal the global dynamics of Hamiltonian systems. We find that the mapping of a slice of the phase space of the HH system obtained using the Poincaré surface of section (PSS) technique [39] and the GALI method leads to the same qualitative features in the identification of regions of regular and chaotic motion. In addition, by recording the central processing unit (CPU) time needed to perform the scan of the PSS using the GALI and the MLE implemented with the MPM, we show that the GALI method is still faster compared to the MLE in performing this task.
@@
The content of the paper is organized as follows. Section 2 contains a brief introduction to the LE and GALI. In Sect. 3, we present the theoretical analysis for the accumulation of numerical errors in the computation of the GALI. In Sec. 4, we examine the conditions of reliability for the computation of the GALI using both the VM and MPM, focusing on regular and chaotic orbits of the HH and the $\beta-$FPUT models. Finally in Sec. 5, we summarize our findings and discuss some open questions. In the appendices, we provide a number of details relevant to this work, including the pseudo code for the computation of GALI using the MPM, initial conditions and additional GALI results for the FPUT model.
2 Lyapunov exponents (LEs) and the Generalized Alignment index (GALI)
@@
In our study we consider autonomous Hamiltonian systems of $N$ degrees of freedom, which are defined by a Hamiltonian function (whose numerical value is typically called the system’s energy) of the form
@@
being the so-called symplectic matrix, where $\bm{0}_{N}$ and $\bm{I}_{N}$ are the $N\times N$ null and identity matrices respectively. In Eq. (2), the over-dot denotes the derivative with respect to time $t$. Considering an initial condition $\bm{x}_{0}$ at time $t_{0}$, the solution of Eq. (2), $\bm{x}(t)$, $t>0$, conserves the value of the system’s energy $\mathcal{H}$ Eq. (1). In addition, we set $t_{0}=0$.
Let us $k$ orthogonal small deviation vectors from a system’s reference orbit $\bm{x}$, namely $\delta\bm{x}_{1}$, $\delta\bm{x}_{2}$, …, $\delta\bm{x}_{k}$, we define the $k$th finite-time LEs
$\lambda_{k}(t)=\frac{1}{t}\ln\frac{\lVert\delta\bm{x}_{k}(t)\rVert}{\lVert\delta\bm{x}_{k}(0)\rVert},$ (4)
@@
$\Lambda_{k}=\lim_{t\rightarrow\infty}\lim_{\delta\bm{x}_{k}(0)\rightarrow 0}\lambda_{k}(t),$ (5)
are called Lyapunov exponents. As $\Lambda_{k}$ is unreachable for practical computation, we mainly rely on the $\lambda_{k}(t)$ at $t\gg 1$ for their approximation. Thus, we use the fact that the finite-time LEs $\lambda_{k}(t)\propto t^{-1}\ln t$ for regular trajectories, and approaches a constant for chaotic trajectories.
Instead of monitoring the rate of growth of each deviation vector, we can also look at the alignments between all of them. Then the GALI of order $k$ ($\mbox{GALI}_{k}$) measures the volume of the generalized parallelepiped whose edges are the corresponding $k$ unit vectors $\widehat{\delta\bm{x}}_{k}(t)=\delta\bm{x}_{k}(t)/\lVert\delta\bm{x}_{k}(t)\rVert$, and is given by
@@
where $\wedge$’ denotes the wedge product of vectors and $t=R\tau$.
The ability of the GALI to distinguish between regular and chaotic orbits was theoretically explained in [14] and has been demonstrated in several applications of the index to various dynamical systems [40, 41, 42]. More specifically, the time evolution of $\mbox{GALI}_{k}$ for a regular orbit is [14, 43, 15]
@@
i.e., the index either remains practically constant for $2\leq k\leq N$, or decays to zero following a power law evolution for $N<k\leq 2N$. On the other hand, for chaotic orbits $\mbox{GALI}_{k}$ decrease exponentially fast to zero with an exponent which depends on the values of the $k$ largest LEs $\lambda_{1}\geq\lambda_{2}\geq\ \ldots\geq\lambda_{k}$ [14]
@@
3 Theoretical estimations of the GALI’s numerical evaluation
@@
Our aim is to investigate the numerical errors arising during the numerical computation of the GALI using the MPM and VM. Before proceeding we outline a number of assumptions aiming to simplify our analysis. In the rest of this section we assume the volume formed by the unit tangent vectors to the reference orbits, i.e. the GALI value at any time of the system’s evolution is bounded by the one of a right angled parallelepiped with side the same unit tangent vectors. It follows that it is enough to focus on the product of the magnitudes of the tangent vectors as it shares the same numerical error with GALI up to a multiplicative factor which depends on the orientations between the different tangent vectors.
@@
Furthermore, given that all the tangent vectors to the reference orbit have initially the same magnitude, and due to the fact that they are evolved independently, we will consider that the numerical error accumulates equally for all the tangent vectors. That is to say, it is sufficient to evaluate the numerical errors of one single vector to deduce the numerical errors during the computation of the volume formed by all the tangent vectors, which is proportional to (or of the same order as) GALI.
@@
In addition, we roughly consider that the origin of the numerical errors stems from three main contributions which are independent for all times: (i) the machine precision, (ii) the magnitude of the tangent vectors at the renormalization time (iii) and the global truncation errors of the numerical integration schemes [29]. It follows that the numerical uncertainty for the $\mbox{GALI}_{k}$ is complex and expressed as a multivariate polynomial of order $k$ whose variables are all factors influencing the numerical evolution of the phase space and tangent vectors. In order to analyse the error meaningfully, we consequently focus only on the leading order terms, because they are simple and depend on the main extensive parameters influencing the evolution of the tangent vectors. We will show that such terms are the primary contributors to the numerical inaccuracies of the computed GALI in short-time scales.
@@
For completeness, we point toward additional notes relevant to this work in the literature. The effects of the global truncation errors of the numerical integration scheme, the length of the initial tangent vectors, the renormalization time and the machine precision on the computation of the MLE are investigated in Ref. [29]. Furthermore, in Sec. 3.5 of Ref. [3] (and references therein), additional sources of numerical errors in the computation of the LEs are studied. For instance, the numerical orthogonalization’s related errors and the statistical errors due to the finite time computations of the LEs. In the case of the GALI method, the effects of the numerical integration techniques during its numerical computation are briefly investigated by Ref. [26].
3.0.1 The variational method

### v2

@@ Abstract
We present a method for the computation of the Generalized Alignment Index (GALI), a fast and effective chaos indicator, using a multi-particle approach that avoids variational equations. We show that this approach is robust and accurate by deriving a leading-order error estimation for both the variational (VM) and the multi-particle (MPM) methods, which we validate by performing extensive numerical simulations on two prototypical models: the two degrees of freedom Hénon-Heiles system and the multidimensional $\beta$-Fermi-Pasta-Ulam-Tsingou chain of oscillators. The dependence of the accuracy of the GALI on control parameters such as the renormalization time, the integration time step and the deviation vector size is studied in detail. We test the MPM implemented with double precision accuracy ($\varepsilon\approx 10^{-16}$) and find that it performs reliably for deviation vector sizes $d_{0}\approx\varepsilon^{1/2}$, renormalization times $\tau\lesssim 1$, and relative energy errors $E_{r}\lesssim\varepsilon^{1/2}$. These results hold for systems with many degrees of freedom and demonstrate that the MPM is a robust and efficient method for studying the chaotic dynamics of Hamiltonian systems. Our work makes it possible to explore chaotic dynamics with the GALI in a vast number of systems by eliminating the need for variational equations.
journal: Communications in Nonlinear Science and Numerical Simulation
@@ 1 Introduction
Revealing the properties of individual or ensembles of orbits is of fundamental interest for nonlinear systems. A key set of tools for this task are chaos indicators, real-valued quantities which exhibit distinct behaviors for chaotic and regular dynamics [1, 2, 3, 4]. Preeminent among these indicators are the Lyapunov exponents (LEs) [5, 6] which quantify chaos [7, 4, 8, 9] and can indicate the existence of first integrals of the dynamics [10, 11] in nonlinear systems, based on the growth rate of small deviations to a given trajectory. Following the LEs, the importance of detecting chaos has led to a plethora of new techniques.These include the fast Lyapunov indicator (FLI) [12, 13], the mean exponential growth of nearby orbits (MEGNO) [14, 15] the smaller (SALI) and generalized (GALI) alignment indices [16, 17, 18], which have found applications in a broad array of fields such as information theory [19], biophysics and materials [20], astrophysics [21, 22, 23], as well as population dynamics and epidemiology [24, 25, 26].
The wide-ranging importance of these indicators has motivated extensive work on how to effectively and efficiently detect chaos, whether through optimizing the numerical procedures [27, 28, 29], choosing appropriate methods of calculation [30, 31], or using short-time computations to determine chaoticity [32]. The GALI method has been particularly successful in reliably and rapidly identifying chaos [17, 33], and is as such worthy of investment into finding ways of optimizing and generalizing its applicability in nonlinear dynamics.
These chaos indicators, including the GALI, rely on the time evolution of deviation (tangent) vectors to a trajectory under investigation. One approach for obtaining these deviation vectors is to generate a reference trajectory along with trajectories starting nearby. Then the deviation vectors are calculated directly by estimating the separation vectors between the nearby trajectories and the reference one – an approach we term the multi-particle method (MPM). Depending on whether the neighboring orbits are kept in the vicinity of, or are driven away from, the system’s reference orbit allows the characterization of the studied trajectory’s long-time behavior. On the other hand, since the nearby trajectories have to be infinitesimally close to the reference one in the phase space, one can generate the dynamics of these deviation vectors by linearizing the system’s equations of motion around the studied trajectory, yielding the so-called variational equations. The set of the equations of motion and the variational equations can therefore be integrated directly. This method is commonly known as the variational method (VM).
Due to its accuracy and efficiency, the VM is typically used for chaos detection [34] and as such has been the preferred approach in the definition and construction of recent chaos indicators. Nevertheless, the VM requires the analytical expression of the equations of motion which must be continuous and differentiable over the considered space-time domain. Such strong requirements cannot always be satisfied, for example in systems in which elements interact following Hertzian contacts [35], non-analytic softening/hardening laws [36, 37] and in systems defined by empirical fields as in complex chemical reactions [38, 39, 40]. Further, even in cases where it is possible for the variational equations to be derived, as models become more complex these equations can be impractical and computationally inefficient to work with (see e.g. [41]).
We note that the MPM framework has been proposed to address the shortcomings of the VM for the computation of the LEs. In this context, several studies on the reliability of the MPM for the calculation of LEs have already been carried out [30, 28, 31] (see also Sect. 3.5 of [4]). In these works it was found that the computation of the LEs can be affected by factors like the size of the initial separation vectors or the renormalization procedures. It is worth noting that arbitrary choices of control parameters for the computation of LEs using the MPM can lead to spurious estimates as observed for example in false positive MLE estimations in planetary motion [30]. These results raise the question of how other chaos indices perform within the MPM framework, an area that remains largely unexplored. In this study, we focus on the GALI method, seeking to resolve several unanswered questions regarding its computation in the context of MPM. For instance, can the GALI be effectively defined within the MPM framework, and how does the evaluated quantity compare to its computation using the VM? How are GALI computations influenced by the MPM? Finally, can GALI, when computed via the MPM, accurately and efficiently detect chaotic behavior?
@@
The aim of our study is to investigate the reliability of the GALI computation through the MPM. To achieve this goal we introduce a definition for the GALI based on the MPM formulation and demonstrate that it is equivalent to using the VM. We then obtain leading order analytical expressions of the numerical accuracy during the computation of the GALI values for both methods, and validate our analytical results by simulating the short-time dynamics of some well-known, prototypical nonlinear Hamiltonian systems: the Hénon-Heiles (HH) [42] and $\beta$-Fermi-Pasta-Ulam-Tsingou ($\beta$-FPUT) [43, 44] models.
We find optimal numerical parameters for computing the GALI with the MPM, and show that this computation enables full-scale nonlinear dynamics investigations at the same accuracy level as using the VM. We also examine the ability of the GALI method to reveal the global dynamics of Hamiltonian systems, once again probing the HH system, recapitulating known results and providing quantitative insights into the chaotic dynamics. These findings show that with well-chosen parameters it is entirely possible to use the MPM to accurately calculate the GALI, opening the way for a large variety of future studies of chaotic dynamics in complex systems.
@@
The content of the paper is organized as follows. Section 2 contains a brief introduction to the LEs and the GALI. In Sect. 3, we present the theoretical analysis for the accumulation of numerical errors in the computation of the GALI. In Sect. 4, we examine the conditions of reliability for the computation of the GALI using both the VM and MPM, focusing on regular and chaotic orbits of the HH and the $\beta$-FPUT models. Finally in Sect. 5, we summarize our findings and discuss some open questions. In the appendices, we provide a number of details relevant to this work, including a pseudo-code for the computation of GALI using the MPM, initial conditions for some orbits considered in the main body of the paper, and additional GALI results for the $\beta$-FPUT model.
2 Lyapunov exponents and the generalized alignment index
@@
In our study we consider autonomous Hamiltonian systems of $N$ degrees of freedom ($N$D), which are defined by a Hamiltonian function (whose numerical value is typically called the system’s energy) of the form
@@
being the so-called symplectic matrix, where $\bm{0}_{N}$ and $\bm{I}_{N}$ are the $N\times N$ null and identity matrices respectively. In Eq. (2), the over-dot denotes the derivative with respect to time $t$. Considering an initial condition $\bm{x}_{0}$ at time $t_{0}$, the solution of Eq. (2), $\bm{x}(t)$, $t>0$, conserves the value of the system’s energy $\mathcal{H}$ [Eq. (1)]. In addition, we set $t_{0}=0$.
An $N$D Hamiltonian system has $2N$ (possibly non-distinct) LEs, ordered as $\Lambda_{1}\geq\Lambda_{2}\geq\cdots\geq\Lambda_{2N}$ In the case of autonomous Hamiltonian systems, these exponents come in pairs of opposite signs, implying that $\sum_{i=1}^{2N}\Lambda_{i}=0$. In addition, due to the presence of an integral of motion (the Hamiltonian function itself), at least two of the LEs are always equal to zero [5, 6, 7, 4, 45]. The most commonly used chaos indicator is the estimation of the maximum LE (MLE) $\Lambda_{1}$, as $\Lambda_{1}>0$ indicates chaos, while $\Lambda_{1}=0$ denotes regular motion. In order to compute the MLE we follow the time evolution of a deviation vector $\delta\bm{x}$ from the system’s reference orbit $\bm{x}$, and define the finite-time MLE
$\lambda_{1}(t)=\frac{1}{t}\ln\frac{\lVert\delta\bm{x}(t)\rVert}{\lVert\delta\bm{x}(0)\rVert},$ (4)
@@
$\Lambda_{1}=\lim_{t\rightarrow\infty}\lim_{\delta\bm{x}(0)\rightarrow 0}\lambda_{1}(t).$ (5)
As $\Lambda_{1}$ is unreachable for practical computations, we mainly rely on the evaluation of $\lambda_{1}(t)$ at $t\gg 1$ for its approximation. Thus, we use the fact that the finite-time MLE evolves as $\lambda_{1}(t)\propto t^{-1}\ln t$ for regular trajectories, and approaches a positive constant value for chaotic trajectories.
A numerical technique for the computation of all LEs, based on the time evolution of many deviation vectors $\delta\bm{x}_{i}$, $i=1,2,\ldots,k$, with $k\leq 2N$, which are kept linearly independent through a Gram-Schmidt orthonormalization procedure is given in [5, 6]. According to this method (which is usually called the ‘standard method’) all other LEs $\Lambda_{2}$, $\Lambda_{3}$ etc., apart from the MLE obtained from (5), are computed as the limits for $t\rightarrow\infty$ and $\delta\bm{x}_{i}(0)\rightarrow 0$ of some appropriate quantities $\lambda_{2}(t)$, $\lambda_{3}(t)$ etc., which are called the finite time LEs (see [6, 7] for more details).
Another approach for identifying the chaoticity of orbits is obtained by monitoring the alignment between $k$ initially orthogonal deviation vectors. Then the GALI of order $k$ ($\mbox{GALI}_{k}$) measures the volume of the generalized parallelepiped whose edges are the corresponding $k$ unit vectors $\widehat{\delta\bm{x}}_{k}(t)=\delta\bm{x}_{k}(t)/\lVert\delta\bm{x}_{k}(t)\rVert$, and is given by
@@
where $\wedge$’ denotes the wedge product of vectors and $t=R\tau$ are the times at which the evolved deviation vectors are again normalized to their initial norm value.
The ability of the GALI method to distinguish between regular and chaotic orbits was theoretically explained in [17] and has been demonstrated in several applications of the index to various dynamical systems (see e.g. [46, 47, 48]). More specifically, the time evolution of the $\mbox{GALI}_{k}$ for a regular orbit is [17, 18]
@@
i.e., the index either remains practically constant for $2\leq k\leq N$, or decays to zero following a power law evolution for $N<k\leq 2N$. On the other hand, for chaotic orbits the $\mbox{GALI}_{k}$ decreases exponentially fast to zero with an exponent which depends on the values of the $k$ largest LEs $\lambda_{1}\geq\lambda_{2}\geq\ \ldots\geq\lambda_{k}$ [17]
@@
3 Theoretical estimations of the generalized alignment index numerical evaluation
@@
Our aim is to investigate the numerical errors arising during the numerical computation of the GALI using the MPM and the VM. Before proceeding we outline a number of assumptions aiming to simplify our analysis. In the rest of this section we assume that the volume formed by the unit tangent vectors to the reference orbit, i.e., the GALI value at any time of the system’s evolution, is bounded by the volume of a right angled parallelepiped with sides the same unit tangent vectors. It follows that it is enough to focus on the product of the magnitudes of the tangent vectors as it shares the same numerical error with the GALI up to a multiplicative factor which depends on the orientations between the different tangent vectors.
@@
Furthermore, given that all the tangent vectors to the reference orbit have initially the same magnitude, and due to the fact that they are evolved independently, we consider that the numerical error accumulates equally for all tangent vectors. That is to say, it is sufficient to evaluate the numerical error of one single vector to deduce the numerical error during the computation of the volume formed by all tangent vectors, which is proportional to (or of the same order as) the GALI’s error.
@@
In addition, we roughly consider that the origin of the numerical errors stems from three main contributions which are independent for all times: (i) the machine precision, (ii) the magnitude of the tangent vectors at the renormalization time, and (iii) the global truncation errors of the numerical integration schemes [31]. Therefore, the numerical error in the computation of the $\mbox{GALI}_{k}$ is intricate and can be represented as a multivariate polynomial of order $k$ with constant coefficients, whose variables correspond to the control parameters above. In order to analyze the error meaningfully, we consequently focus only on the leading order terms, because they are simple and depend on the main extensive parameters influencing the numerical evolution of the phase space orbits and their associated tangent vectors. We will show that such terms are the primary contributors to the numerical inaccuracies of the computed GALI in short-time scales.
@@
For completeness, we point toward additional notes relevant to this work in the literature. The effects of the global truncation errors of the numerical integration scheme, the length of the initial tangent vectors, the renormalization time and the machine precision on the computation of the MLE are investigated in Ref. [31]. Furthermore, in Sect. 3.5 of Ref. [4] (and references therein), additional sources of numerical errors in the computation of the LEs are studied. For instance, the numerical orthogonalization’s related errors and the statistical errors due to the finite time computations of the LEs. In the case of the GALI method, the effects of the numerical integration techniques during its numerical computation are briefly investigated in Ref. [28].
3.1 The variational method